The market determined price in a perfectly competitive industry is P = Rs. 10. Suppose that the total cost equation of an individual firm in the industry is given by the expression 

TC 1000+2Q+0.01Q2 

a) What is the firm’s profit-maximizing output level and profit? Is this profit normal profit or supper normal profit? Justify your answer (10 marks)

b) At profit maximizing level what is firm total cost, total revenue and marginal cost (10 Marks)

c) Why does a competitive firm is considered as a price taker and Monopoly firm as a price maker (05 Marks)


1
Expert's answer
2021-08-20T08:51:04-0400

Solution:

a.). A firm in a perfectly competitive industry profit-maximizing output level is determined at the point where MR = MC, or where P = MC.

P = 10

TC = 1000 + 2Q + 0.01Q2


MC = "\frac{\partial TC} {\partial Q}" = 2 + 0.02Q


Set P = MC:

10 = 2 + 0.02Q

10 – 2 = 0.02Q

8 = 0.02Q

Q = 400

The firm’s profit-maximizing level of output = 400


Profit = TR – TC

TR = P "\times" Q = 10 "\times" 400 = 4,000

TC = 1000 + 2(400) + 0.01(4002) = 1000 + 800 + 1600 = 3,400

Profit = 4,000 – 3,400 = 600


The firm’s profit is Rs. 400, which is a supernormal profit. This is because the firm’s total revenue exceeds the total costs of production with a huge margin, that is the firm is earning a profit above and beyond the level of normal profit.


b.). At profit-maximizing level what is the firm total cost, total revenue, and marginal cost:

At profit maximizing level: Q = 400 and P = 10

Total cost = 1000 + 2(400) + 0.01(4002) = 1000 + 800 + 1600 = 3,400

TC = 3,400

 

Total revenue = P "\times" Q = 10 "\times" 400 = 4,000

TR = 4,000

 

Marginal cost = 2 + 0.02Q = 2 + 0.02(400) = 2 + 8 = 10

MC = 10


c.). A competitive firm is considered a price taker since the firm cannot influence the price in the market and therefore, it must take the market price as given or what other firms in the market are charging. Similarly, the pressure from competing firms forces the firm to take the prevailing market equilibrium price. Should the firm increase its prices, it would lose its sales revenue to its competitors.

 

A monopoly firm is considered a price maker since it has the power and influence over the market prices due to the lack of perfect substitutes for its products. A monopoly sells unique products that have no close substitutes hence the power to charge any price. Also, due to the absence of competition from other firms, a monopoly firm can charge any price for its products without losing its customers or sales revenue.